# Robust Fast Adaptation from Adversarially Explicit Task Distribution Generation

### v1

@@ 5. Experiments
Previous sections recast the adversarially task robust meta-learning to a Stackelberg game, specify the equilibrium, and analyze theoretical properties in distribution generation. This section focuses on the evaluation, and baselines constructed from typical risk minimization principles are reported in Table (3). These include vanilla MAML (Finn et al., 2017), DRO-MAML (Collins et al., 2020), TR-MAML (Collins et al., 2020), DR-MAML (Wang et al., 2023c), and AR-MAML (ours).
@@ 5.1. Benchmarks
We consider the few-shot synthetic regression, system identification, and meta reinforcement learning to test fast adaptation robustness with typical baselines. Notably, the task is specified by the generated task identifiers as shown in Fig. (3).
@@
We refer the reader to Appendix (I) for set-ups, hyper-parameter configurations and additional experimental results.
@@ 5.2. Empirical Result Analysis
Overall Performance: Table (1) shows that AR-MAML mostly outperforms others in the adversarial distribution, seldom sacrificing performance in the initial distribution. Similar to observations in (Wang et al., 2023c), task distributionally robust optimization methods, like DR-MAML and DRO-MAML, not only retain robustness advantage on shifted distribution but also sometimes boost average performance on the initial distribution. Cases with two types of initial task distributions (Uniform/Normal) come to similar conclusions on average and $\text{CVaR}_{\alpha}$ performance. Figs. (4)/(5) show the meta reinforcement learning results for Point Robot and Ant Pos navigation tasks. AR-MAML exhibits similar superiority on both continuous control benchmarks compared to baselines.
@@
Multiple Tail Risk Robustness: Note that CVaR metrics imply the model’s robustness under the subpopulation shift. Fig. (6) reports $\text{CVaR}_{\alpha}$ values with various confidence values on pendulum system identification. The AR-MAML’s merits in handling the proportional worst cases are consistent across diverse levels. We also illustrate and include these statistics on other benchmarks in Appendix (J). Moreover, as suggested in (Taori et al., 2020), a robust learner seldom encounters a performance gap between a standard (initial) test set and a test set with a distribution shift (adversarial). Fig. (7) validates the meta-learners’ robustness on sinusoid regression, where AR-MAML’s results are more proximal to the $y=x$ line than other baselines.
@@
Random Perturbation Robustness: We also test meta-learners’ robustness to random noise from the support dataset. To do so, we take sinusoid regression and inject random noise into the support set, i.e., the noise is drawn from a Gaussian distribution $\mathcal{N}(0,0.1^{2})$ and added to the output $y$. Fig. (8) illustrates that AR-MAML’s performance degradation is somewhat less than others on the adversarial distribution. The noise exhibits similar effects on AR-MAML and DR-MAML on the initial distribution, harming performance severely. AR-MAML and DR-MAML still exhibit lower MSEs than other baselines for all cases. This indicates the adversarial training mechanism can also bring more robustness to challenging test scenarios with random noise.
@@ 5.3. Task Structure Analysis
Explicit Task Distribution: As displayed in Fig. (9), our approach enables the discovery of explicit task structures regarding problem-solving. The general learned patterns seem to be regardless of the initial task distributions. In sinusoid regression, more probability mass is allocated in the region with $[3.0,5.0]\times[0.0,1.0]$, which reveals more difficulties in adaptation with larger amplitude descriptors. For the Pendulum, the distribution adversary assigns less probability mass to two corner regions, implying that the combination of higher masses and longer pendulums or lower masses and shorter pendulums is easier to predict. Similar phenomena are observed in mass combinations of Acrobat systems. Consistently, the existence of constraint decreases all task distribution entropies to a certain level, which we report in Appendix (I).
Initial Task Distributions’ Influence on Structures: Comparing the top and the bottom of Fig. (9), we notice that the uniform and the normal initial distribution results in similar patterns after normalizing flows’ transformations on separate benchmarks. The normal initial distribution can be transformed into smooth ones and captures high-density regions around centroids.
@@ 5.4. Other Investigations
Impacts of Shift Distribution Constraints: Our studied framework allows the task distribution to shift at a certain level. In Eq. (6), larger $\lambda$ values tend to cause the generated distribution to collapse into the initial distribution. Consequently, we empirically test the naive and severe adversarial training, e.g., setting $\lambda=\{0.0,0.1,0.2\}$ on sinusoid regression. As displayed in Fig. (10), the generated distribution with $\lambda=0.0$ suffers from severe mode collapse, merely covering diagonal regions in the task space. Such a curse is alleviated with increasing $\lambda$ values. In Fig. (11), the meta learner, after heavy distribution shifts, catastrophically fails to generalize well in the initial distribution, illustrating higher adaptation risks in $\lambda=0.0$.
Figure 10. Adversarial Task Probability Distribution on Sinusoid Regression with Various Lagrange Multiplier $\lambda$.
@@
Compatibility with Other Meta-learning Methods: Besides the AR-MAML, we also check the effect of adversarially task robust training with other meta-learning methods. Here, AR-CNP in Example (2) is employed in the evaluation. Take the sinusoid regression as an example. Table (2) observes comparable performance between AR-CNP and DR-CNP on the initial task distribution, while results on the adversarial task distribution uncover a significant advantage over others, particularly on robustness metrics, namely $\text{CVaR}_{\alpha}$ values.
@@ Acknowledgements.
This work is funded by National Natural Science Foundation of China (NSFC) with the Number # 62306326. And we thank Yuhang Jiang, Chen Chen, and Daming Shi for helpful discussions.
@@ A.2. Players’ Order in Decision-making
The Remark (3) indicates the game is asymmetric w.r.t. the players’ decision-making order. This work abstracts the decision-making process for the distribution adversary and the meta-learner in a sequential game. Hence, the order inevitably influences the solution concept. Actually, we take MAML as an example to implement the adversarially task robust meta-learning framework. As noted in Algorithm (1), updating the distribution adversary’s parameter requires the evaluation of task-specific model parameters, which are not available in the initialization. The traditional game theory no longer applies to our studied case, and simple counter-examples such as rock/paper/scissors show no equilibrium in practice. Hence, we pre-determine the order of decision-makers for implementation, as reported in the main paper.
@@ A.5. Related Work Summary
As far as we know, robust fast adaptation is an underexplored research issue in the field. In principle, we include the latest SOTA (Wang et al., 2023c; Collins et al., 2020) and typical SOTA (Sagawa et al., 2019) methods to compare in this work as illustrated in Table (3). Note that in (Wang et al., 2023c), the idea of increasing the robustness is to introduce the task selector for risky tasks in implementations. Though there are some curriculum methods (Yao et al., 2021b; Dennis et al., 2020; Parker-Holder et al., 2022) to reschedule the task sampling probability along the learning, our setup focuses on (i) the semantics in the task identifiers to explicitly uncover the task structure from a Stackelberg game and (ii) the robust solution search under a distribution shift constraint. That is, ours requires capturing the differentiation information from the task identifiers of tasks, such as masses and lengths in the pendulum system identification. Ours is agnostic to meta-learning methods and avoids task distribution mode collapse, which might happen in adversarial training without constraints. These configuration and optimization differences drive us to include the risk minimization principle in the comparisons.
@@ Appendix D Theoretical Understanding of Generating Task Distribution
Tolerant Region in distribution shifts. Next, we interpret the above-induced optimization objective. As exhibited in Eq. (24), the condition constrains the generative task distribution $p_{\bm{\phi}}(\tau)$ within a neighborhood of the initial task distribution $p_{0}(\tau)$. This can be viewed as the tolerant region of task distribution shifts, which means larger $\delta$ allows more severe distribution shifts concerning the initial task distribution. As for the role of the distribution adversary, it attempts to seek the task distribution that can handle the strongest task distribution shift under the allowed region, as displayed in Fig. (12).
@@ E.1. Existence of Stackelberg Equilibrium
Note that for the nonconvex-nonconcave min-max optimization problem, there is no general guarantee for the existence of saddle points. The solution concept in the Definition (1) requires no convexity w.r.t. the optimization objective and provides a weaker equilibrium, also referred to as the global minimax point (Jin et al., 2020).
With the Assumption (1) and the Definition (1), the global Stackelberg equilibrium always exists for the proposed min-max optimization $\min_{\bm{\theta}\in\bm{\Theta}}\max_{\bm{\phi}\in\bm{\Phi}}\mathcal{J}(\bm{\theta},\bm{\phi})$. However, its exact search is NP-hard, and the stochastic optimization practically leads to the local Stackelberg equilibrium in the Definition (4).
@@ E.2. Convergence Guarantee
Theorem 1 (Convergence Guarantee) Suppose the Assumption (1) and the condition of the (local) Stackelberg equilibrium $\Delta(\mathbf{A},\mathbf{B},\mathbf{C},\gamma_{1},\gamma_{2}):=\max\Big{\{}(1-\gamma_{1}\sigma_{\text{min}})^{2}(1+\gamma_{2}^{2}L_{\text{max}}^{2}),\Big{|}|\gamma_{1}^{2}-2\gamma_{1}\gamma_{2}+\gamma_{1}^{2}\gamma_{2}^{2}L_{\text{max}}^{2}|L_{\text{max}}^{2}+(1+\gamma_{2}L_{\text{max}})^{2}-2\gamma_{1}\gamma_{2}^{2}\sigma_{\text{min}}(\mathbf{B}^{T}\mathbf{B}\mathbf{C})\Big{|}\Big{\}}<\frac{1}{2}$ are satisfied, where $\sigma_{\text{min}}(\cdot)$ denotes the smallest eigenvalues of the corresponding matrix. Then we can have the following statements:
@@
when using the alternating GDA for solving the adpatively robust meta-learning. Fig. (13) illustrates the iteration rules and steps.
@@ Appendix F Generalization Bound
Note that the task distribution is adaptive and learnable, and we take interest in the generalization in the context of generative task distributions. It is challenging to perform direct analysis. Hence, we propose to exploit the importance of the weighting trick. To do so, we first recap the reweighted generalization bound from (Cortes et al., 2010) as Lemma (1). The sketch of proofs mainly consists of the importance-weighted generalization bound and estimates of the importance weights’ range.
@@ Appendix H Evolution of Entropies in Task Distributions
One of the primary benefits of employing normalizing flows in generating task distributions is the tractable likelihood, leaving it possible to analyze the entropy of task distributions. Note that Remark (1) in the main paper characterizes the way of computing the generative task distribution entropy.
@@
This completes the proof of Remark (1).
@@ I.1. Meta Learning Benchmarks
Sinusoid Regression. In Sinusoid regression, each task is equivalent to mapping the input to the output of a sine wave. Here, the task identifiers are the amplitude $a$ and the phase $b$. Data points in regression are collected in the way: $10$ data points are uniformly sampled from the interval $[-5.0,5.0]$, coupled with the output $y=a\sin(x+b)$. These data points are divided into the support dataset ($5$-shot) and the query dataset. As for the range of task identifiers and the types of initial distributions, please refer to Table (4). In meta-testing phases, we randomly sample $500$ tasks from the initial task distribution and the generated task distribution to evaluate the performance, and this results in Table (1).
@@
System Identification. In the Acrobot System, angles and angular velocities characterize the state of an environment as $[\bm{\theta}_{1},\bm{\theta}_{1}^{\prime},\bm{\theta}_{2},\bm{\theta}_{2}^{\prime}]$. The goal is to identify the system dynamics, namely the transited state, after selecting Torque action from $\{-1,0,+1\}$. In the Pendulum System, environment information can be found in the OpenAI gym. In detail, the observation is $(\cos\bm{\theta},\sin\bm{\theta},\bm{\theta}^{\prime})$ with $\bm{\theta}\in[-\pi,\pi]$, and the continuous action range is $[-2.0,2.0]$. The torque is executed on the pendulum body, and the goal is to predict the dynamics given the observation and action. For both dynamical systems, we use a random policy as the controller to interact with the environment to collect transition samples. For the few-shot purpose, we sample $200$ transitions from each Markov Decision Process as one batch and randomly split them into the support dataset ($10$-shot transitions) and the query dataset. In meta training, the meta-batch in iteration is $16$, and the maximum iteration number is $500$. In meta-testing phases, we randomly sample $100$ tasks respectively from the initial and generated task distributions with each task one batch in evaluation, which results in Table (1).
@@
Meta Learning Continuous Control. We consider the navigation task to examine the performance of methods in reinforcement learning. The mission is to guide the robot, e.g., the point robot and the Ant from Mujoco, to move towards the target goal step by step. Hence, the task identifier is the goal location $(x_{1},x_{2})$. The agent performs $20$ episode explorations to identify the environment and enable inner policy gradient updates as fast adaptation. The environment information, such as transitions and rewards, is accessible at (https://github.com/lmzintgraf/cavia/tree/master/rl/envs). Particularly, for the task distribution like $\mathcal{U}([-0.5,0.5]\times[-0.5,0.5])$ in the point robot, we set the dark area in Fig. (3c) as the sparse reward region, discounting the step-wise reward by $0.6$ to aera $[-0.25,-0.5]\times[0.25,0.5]$ and $0.4$ to aera $[0.25,0.5]\times[-0.25,-0.5]$. In meta training, the meta-batch in the iteration is 20 for the point robot, and 40 for Ant, and the model is trained for up to 500 meta-iterations. In meta-testing, we randomly sampled 100 MDPs, specifically from the initial and generated task distributions. For each MDP, we run 20 episodes as the support dataset and compute the cumulated returns after fast adaptation.

### v3

@@ 5. Experiments
Previous sections recast the adversarially task robust meta-learning to a Stackelberg game, specify the equilibrium, and analyze theoretical properties in distribution generation. This section focuses on the evaluation, and baselines constructed from typical risk minimization principles are reported in Table 3. These include vanilla MAML (Finn et al., 2017), DRO-MAML (Sagawa et al., 2019), TR-MAML (Collins et al., 2020), DR-MAML (Wang et al., 2023c), and AR-MAML (ours).
@@ 5.1. Benchmarks
We consider the few-shot synthetic regression, system identification, and meta reinforcement learning to test fast adaptation robustness with typical baselines. Notably, the task is specified by the generated task identifiers as shown in Figure 3.
@@
We refer the reader to Appendix I for set-ups, hyper-parameter configurations and additional experimental results.
@@ 5.2. Empirical Result Analysis
Overall Performance: Table 1 shows that AR-MAML mostly outperforms others in the adversarial distribution, seldom sacrificing performance in the initial distribution. Similar to observations in (Wang et al., 2023c), task distributionally robust optimization methods, like DR-MAML and DRO-MAML, not only retain robustness advantage on shifted distribution but also sometimes boost average performance on the initial distribution. Cases with two types of initial task distributions (Uniform/Normal) come to similar conclusions on average and $\text{CVaR}_{\alpha}$ performance. Figures 4/5 show the meta reinforcement learning results for Point Robot and Ant Pos navigation tasks. AR-MAML exhibits similar superiority on both continuous control benchmarks compared to baselines.
@@
Multiple Tail Risk Robustness: Note that CVaR metrics imply the model’s robustness under the subpopulation shift. Figure 6 reports $\text{CVaR}_{\alpha}$ values with various confidence values on pendulum system identification. The AR-MAML’s merits in handling the proportional worst cases are consistent across diverse levels. We also illustrate and include these statistics on other benchmarks in Appendix J. Moreover, as suggested in (Taori et al., 2020), a robust learner seldom encounters a performance gap between a standard (initial) test set and a test set with a distribution shift (adversarial). Figure 7 validates the meta-learners’ robustness on sinusoid regression, where AR-MAML’s results are more proximal to the $y=x$ line than other baselines.
@@
Random Perturbation Robustness: We also test meta-learners’ robustness to random noise from the support dataset. To do so, we take sinusoid regression and inject random noise into the support set, i.e., the noise is drawn from a Gaussian distribution $\mathcal{N}(0,0.1^{2})$ and added to the output $y$. Figure 8 illustrates that AR-MAML’s performance degradation is somewhat less than others on the adversarial distribution. The noise exhibits similar effects on AR-MAML and DR-MAML on the initial distribution, harming performance severely. AR-MAML and DR-MAML still exhibit lower MSEs than other baselines for all cases. This indicates the adversarial training mechanism can also bring more robustness to challenging test scenarios with random noise.
@@ 5.3. Task Structure Analysis
Explicit Task Distribution: As displayed in Figure 9, our approach enables the discovery of explicit task structures regarding problem-solving. The general learned patterns seem to be regardless of the initial task distributions. In sinusoid regression, more probability mass is allocated in the region with $[3.0,5.0]\times[0.0,1.0]$, which reveals more difficulties in adaptation with larger amplitude descriptors. For the Pendulum, the distribution adversary assigns less probability mass to two corner regions, implying that the combination of higher masses and longer pendulums or lower masses and shorter pendulums is easier to predict. Similar phenomena are observed in mass combinations of Acrobat systems. Consistently, the existence of constraint decreases all task distribution entropies to a certain level, which we report in Appendix I. Though such a decrease brings more concentration on some task subsets, AR-MAML still probably fails to cover other challenging combinations in mode collapse.
Initial Task Distributions’ Influence on Structures: Comparing the top and the bottom of Figure 9, we notice that the uniform and the normal initial distribution results in similar patterns after normalizing flows’ transformations on separate benchmarks. The normal initial distribution can be transformed into smooth ones and captures high-density regions around centroids.
@@ 5.4. Other Investigations
Impacts of Shift Distribution Constraints: Our studied framework allows the task distribution to shift at a certain level. In Eq. (6), larger $\lambda$ values tend to cause the generated distribution to collapse into the initial distribution. Consequently, we empirically test the naive and severe adversarial training, e.g., setting $\lambda=\{0.0,0.1,0.2\}$ on sinusoid regression. As displayed in Figure 10, the generated distribution with $\lambda=0.0$ suffers from severe mode collapse, merely covering diagonal regions in the task space. Such a curse is alleviated with increasing $\lambda$ values. In Figure 11, the meta learner, after heavy distribution shifts, catastrophically fails to generalize well in the initial distribution, illustrating higher adaptation risks in $\lambda=0.0$.
Figure 10. Adversarial Task Probability Distribution on Sinusoid Regression with Various Lagrange Multipliers $\lambda$.
@@
Compatibility with Other Meta-learning Methods: Besides the AR-MAML, we also check the effect of adversarially task robust training with other meta-learning methods. Here, AR-CNP in Example 2 is employed in the evaluation. Take the sinusoid regression as an example. Table 2 observes comparable performance between AR-CNP and DR-CNP on the initial task distribution, while results on the adversarial task distribution uncover a significant advantage over others, particularly on robustness metrics, namely $\text{CVaR}_{\alpha}$ values.
@@ Acknowledgements.
This work is funded by National Natural Science Foundation of China (NSFC) with the Number # 62306326. And we thank Dong Liang, Yuhang Jiang, Chen Chen, Daming Shi, and other reviewers for helpful discussions.
@@ A.2. Players’ Order in Decision-making
The Remark 3 indicates the game is asymmetric w.r.t. the players’ decision-making order. This work abstracts the decision-making process for the distribution adversary and the meta-learner in a sequential game. Hence, the order inevitably influences the solution concept. Actually, we take MAML as an example to implement the adversarially task robust meta-learning framework. As noted in Algorithm 1, updating the distribution adversary’s parameter requires the evaluation of task-specific model parameters, which are not available in the initialization. The traditional game theory no longer applies to our studied case, and simple counter-examples such as rock/paper/scissors show no equilibrium in practice. Hence, we pre-determine the order of decision-makers for implementation, as reported in the main paper.
@@ A.5. Related Work Summary
As far as we know, robust fast adaptation is an underexplored research issue in the field. In principle, we include the latest SOTA (Wang et al., 2023c; Collins et al., 2020) and typical SOTA (Sagawa et al., 2019) methods to compare in this work as illustrated in Table 3. Note that in (Wang et al., 2023c), the idea of increasing the robustness is to introduce the task selector for risky tasks in implementations. Though there are some curriculum methods (Yao et al., 2021b; Dennis et al., 2020; Parker-Holder et al., 2022) to reschedule the task sampling probability along the learning, our setup focuses on (i) the semantics in the task identifiers to explicitly uncover the task structure from a Stackelberg game and (ii) the robust solution search under a distribution shift constraint. That is, ours requires capturing the differentiation information from the task identifiers of tasks, such as masses and lengths in the pendulum system identification. Ours is agnostic to meta-learning methods and avoids task distribution mode collapse, which might happen in adversarial training without constraints. These configuration and optimization differences drive us to include the risk minimization principle in the comparisons.
@@ Appendix D Theoretical Understanding of Generating Task Distribution
Tolerant Region in distribution shifts. Next, we interpret the above-induced optimization objective. As exhibited in Eq. (24), the condition constrains the generative task distribution $p_{\bm{\phi}}(\tau)$ within a neighborhood of the initial task distribution $p_{0}(\tau)$. This can be viewed as the tolerant region of task distribution shifts, which means larger $\delta$ allows more severe distribution shifts concerning the initial task distribution. As for the role of the distribution adversary, it attempts to seek the task distribution that can handle the strongest task distribution shift under the allowed region, as displayed in Figure 12.
@@ E.1. Existence of Stackelberg Equilibrium
Note that for the nonconvex-nonconcave min-max optimization problem, there is no general guarantee for the existence of saddle points. The solution concept in the Definition 1 requires no convexity w.r.t. the optimization objective and provides a weaker equilibrium, also referred to as the global minimax point (Jin et al., 2020).
With the Assumption 1 and the Definition 1, the global Stackelberg equilibrium always exists for the proposed min-max optimization $\min_{\bm{\theta}\in\bm{\Theta}}\max_{\bm{\phi}\in\bm{\Phi}}\mathcal{J}(\bm{\theta},\bm{\phi})$. However, its exact search is NP-hard, and the stochastic optimization practically leads to the local Stackelberg equilibrium in the Definition 4.
@@ E.2. Convergence Guarantee
Theorem 1 (Convergence Guarantee) Suppose the Assumption 1 and the condition of the (local) Stackelberg equilibrium $\Delta(\mathbf{A},\mathbf{B},\mathbf{C},\gamma_{1},\gamma_{2}):=\max\Big{\{}(1-\gamma_{1}\sigma_{\text{min}})^{2}(1+\gamma_{2}^{2}L_{\text{max}}^{2}),\Big{|}|\gamma_{1}^{2}-2\gamma_{1}\gamma_{2}+\gamma_{1}^{2}\gamma_{2}^{2}L_{\text{max}}^{2}|L_{\text{max}}^{2}+(1+\gamma_{2}L_{\text{max}})^{2}-2\gamma_{1}\gamma_{2}^{2}\sigma_{\text{min}}(\mathbf{B}^{T}\mathbf{B}\mathbf{C})\Big{|}\Big{\}}<\frac{1}{2}$ are satisfied, where $\sigma_{\text{min}}(\cdot)$ denotes the smallest eigenvalues of the corresponding matrix. Then we can have the following statements:
@@
when using the alternating GDA for solving the adaptively robust meta-learning. Figure 13 illustrates the iteration rules and steps.
@@ Appendix F Generalization Bound
Note that the task distribution is adaptive and learnable, and we take interest in the generalization in the context of generative task distributions. It is challenging to perform direct analysis. Hence, we propose to exploit the importance of the weighting trick. To do so, we first recap the reweighted generalization bound from (Cortes et al., 2010) as Lemma 1. The sketch of proofs mainly consists of the importance-weighted generalization bound and estimates of the importance weights’ range.
@@ Appendix H Evolution of Entropies in Task Distributions
One of the primary benefits of employing normalizing flows in generating task distributions is the tractable likelihood, leaving it possible to analyze the entropy of task distributions. Note that Remark 1 in the main paper characterizes the way of computing the generative task distribution entropy.
@@
This completes the proof of Remark 1.
@@ I.1. Meta Learning Benchmarks
Sinusoid Regression. In sinusoid regression, each task is equivalent to mapping the input to the output of a sine wave. Here, the task identifiers are the amplitude $a$ and the phase $b$. Data points in regression are collected in the way: $10$ data points are uniformly sampled from the interval $[-5.0,5.0]$, coupled with the output $y=a\sin(x-b)$. These data points are divided into the support dataset ($5$-shot) and the query dataset. As for the range of task identifiers and the types of initial distributions, please refer to Table 4. In meta-testing phases, we randomly sample $500$ tasks from the initial task distribution and the generated task distribution to evaluate the performance, and this results in Table 1.
@@
System Identification. In the Acrobot System, angles and angular velocities characterize the state of an environment as $[\bm{\theta}_{1},\bm{\theta}_{1}^{\prime},\bm{\theta}_{2},\bm{\theta}_{2}^{\prime}]$. The goal is to identify the system dynamics, namely the transited state, after selecting Torque action from $\{-1,0,+1\}$. In the Pendulum System, environment information can be found in the OpenAI gym. In detail, the observation is $(\cos\bm{\theta},\sin\bm{\theta},\bm{\theta}^{\prime})$ with $\bm{\theta}\in[-\pi,\pi]$, and the continuous action range is $[-2.0,2.0]$. The torque is executed on the pendulum body, and the goal is to predict the dynamics given the observation and action. For both dynamical systems, we use a random policy as the controller to interact with the environment to collect transition samples. For the few-shot purpose, we sample $200$ transitions from each Markov Decision Process as one batch and randomly split them into the support dataset ($10$-shot transitions) and the query dataset. In meta training, the meta-batch in iteration is $16$, and the maximum iteration number is $500$. In meta-testing phases, we randomly sample $100$ tasks respectively from the initial and generated task distributions with each task one batch in evaluation, which results in Table 1.
@@
Meta Learning Continuous Control. The environment of reinforcement learning is mainly treated as a Markov decision process (Mao et al., 2024, 2023). And the meta RL is about the distribution over environments. We consider the navigation task to examine the performance of methods in reinforcement learning. The mission is to guide the robot, e.g., the point robot and the Ant from Mujoco, to move towards the target goal step by step. Hence, the task identifier is the goal location $(x_{1},x_{2})$. The agent performs $20$ episode explorations to identify the environment and enable inner policy gradient updates as fast adaptation. The environment information, such as transitions and rewards, is accessible at (https://github.com/lmzintgraf/cavia/tree/master/rl/envs). Particularly, for the task distribution like $\mathcal{U}([-0.5,0.5]\times[-0.5,0.5])$ in the point robot, we set the dark area in Figure 3c as the sparse reward region, discounting the step-wise reward by $0.6$ to area $[-0.25,-0.5]\times[0.25,0.5]$ and $0.4$ to aera $[0.25,0.5]\times[-0.25,-0.5]$. In meta training, the meta-batch in the iteration is 20 for the point robot, and 40 for Ant, and the model is trained for up to 500 meta-iterations. In meta-testing, we randomly sampled 100 MDPs, specifically from the initial and generated task distributions. For each MDP, we run 20 episodes as the support dataset and compute the accumulated returns after fast adaptation.